Omar has $4.65 in dimes and quarters. The number of dimes is 6 less than the number of quarters. How many of each coin does he have?
Answer by ikleyn(52797)

.

Add 6 dimes to the collection (mentally).


Then the collection will have equal number of dimes and quarters and be worth 465 + 6*10 = 525 cents.


Hence, you can group all coins in this new collection in pairs, 1 dime + 1 quarters in each pair.


Each pair is worth 25+10 = 35 cents, and the number of these pairs is  525/35 = 15.


Hence, the original collection had 15 quarters and 15-6 = 9 dimes.    ANSWER

Solved.